OpenStudy (anonymous):

A ladder of length L = 2.8 m and mass m = 19 kg rests on a floor with coefficient of static friction μs = 0.48. Assume the wall is frictionless. 1)What is the normal force the floor exerts on the ladder? 2)What is the minimum angle the ladder must make with the floor to not slip? A ladder of length L = 2.8 m and mass m = 19 kg rests on a floor with coefficient of static friction μs = 0.48. Assume the wall is frictionless. 1)What is the normal force the floor exerts on the ladder? 2)What is the minimum angle the ladder must make with the floor to not slip? @Physics

OpenStudy (shamim):

the weight of the lader is W=mg=19*9.8=186.2N

OpenStudy (shamim):

we know frm newtons third law of motion, normal force exerted by the floor on the ladder = the weight of the ladder=186.2N

OpenStudy (shamim):

if u need more explanation plz ask me

OpenStudy (shamim):

u know every action has an equal n opposite reaction. its newtons 2nd law of motion

OpenStudy (shamim):

so the weight of the ladder W is a force exerted by the ladder on the floor is equal n opposite to the force exerted by the floor on the ladder. the force on the ladder is called reaction force R. so W=R

OpenStudy (shamim):

i talked abt question 1 above

OpenStudy (shamim):

now question2

OpenStudy (shamim):

we know limiting static frictional force \[F=\mu _{s}R=0.48 \times mg=0.48 \times 19 \times 9.8=89.38N\]

OpenStudy (shamim):

this limiting(maximum) static frictional force will b equal n opposite to the applied(external) force.

OpenStudy (shamim):

i think applied force will come frm weight W. bt hw. i dont know. still thinking

OpenStudy (anonymous):

question 1: normal force is n=mgsina*sina question 2: f=cn(c=coefficient of friction) mgsinacosa=c*mgsina*sina cosa=csina cota=c cota=0.48

OpenStudy (vincent-lyon.fr):

Sorry, but this is not correct. You cannot "propagate" a force like you are doing with weight in your diagram. Actually it is way much simpler: At equilibrium net force equals zero. Then on the vertical, weight and normal reaction by the ground have to cancel out. Hence N ground = W = mg

OpenStudy (vincent-lyon.fr):

Use torque = 0 to find reaction by wall, then you will be able to find friction exerted by the ground.